OpenStudy (magbak):

I will award medal !!!!!!!!!!!! Please Help!!!!!!!! Using complete sentences, explain how you would factor completely x^9 –27.

OpenStudy (magbak):

@jim_thompson5910 @whpalmer4

OpenStudy (ivettef365):

that is formula (a-b) ( a^2 + ab + b^2)

OpenStudy (uri):

Oh yes :D

OpenStudy (anonymous):

The formula for the difference of two cubes is (x - y)(x + xy + y^2) the cube root of x^9 = x^3. the cube root of 27 = 3 x = x^3 y = 3 substitute into the equation (x^3 - 3)(x^6 + 3x^3 + 9)

OpenStudy (magbak):

Can I get a explanation please.

OpenStudy (anonymous):

I explained it :/ Listen you just have to learn the formula :) and only you can understand

OpenStudy (ivettef365):

look at the formula (a-b)(a^2 + ab + b^2) then you have x^9 - 27 x^9 = (x^3)(x^3)(x^3) 27 = (3)(3)(3) so a = x^3 and b = 3

OpenStudy (ivettef365):

did you understand ?

OpenStudy (magbak):

yes thank you

OpenStudy (ivettef365):

yw :)